# The myth of declining competitive balance in the UEFA Champions League group stage

### v4

@@ Abstract
The degree of competitive balance is a crucial issue in the analysis of sports competitions. According to previous studies, competitive balance has significantly declined in the UEFA Champions League group stage over the recent decades. Our paper introduces six alternative indices in order to explore the robustness of these results. The ex ante measures are based on Elo ratings, while our ex post measures compare the group ranking to reasonable benchmarks. We find no evidence of any trend in the competitive balance of the UEFA Champions League group stage between the 2003/04 and 2023/24 seasons.
@@ 1 Introduction
One of the most prestigious football tournaments, the UEFA Champions League, has been organised in the same format over 21 years, which has fundamentally been changed from the 2024/25 season. Commenting on this decision, the president of the Union of European Football Associations (UEFA), Aleksander Čeferin, said (UEFA,, 2022): “We are convinced that the format chosen strikes the right balance and that it will improve the competitive balance and generate solid revenues that can be distributed to clubs, leagues and into grassroots football across our continent while increasing the appeal and popularity of our club competitions.” Indeed, Gyimesi, (2024) has demonstrated that the new format can be successful in improving competitive balance.
Recent studies have observed a robust decline in both the ex ante (Triguero-Ruiz and Avila-Cano,, 2023) and ex post (Ramchandani et al.,, 2023; Triguero-Ruiz and Avila-Cano,, 2023) competitive balance of the UEFA Champions League group stage over the last decades, which might explain the reform. Our paper aims to check the sensitivity of these results by introducing and explaining alternative indicators of competitive balance. Naturally, the proposed measures may have their own disadvantages and shortcomings, and it is not claimed that they should be preferred to existing indices. However, it would be unreasonable to dismiss them for the sake of the standard measures of competitive balance since “there are no easy formulas for competitive balance success” (Zimbalist,, 2002, p. 120). We think it is the responsibility of the academic community to confirm or refute previous results on competitive balance by considering as many measures as possible, which will be done in the following.
@@
In contrast to the conclusions of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023), no evidence is found of any decline in the competitive balance of the UEFA Champions League group stage in the recent decades, between the 2003/04 and 2023/24 seasons. Consequently, if UEFA has chosen the barely used incomplete round-robin design of the UEFA Champions League with its inherent risks (Devriesere et al.,, 2024) because of the worsening trend in competitive balance, the decision-makers might have been misled.
@@
The paper has the following structure. Section 2 provides a concise literature review. Section 3 summarises some possible limitations of previous studies. Our measures of competitive balance are introduced in Section 4 together with a description of the underlying data. Section 5 provides the results and Section 6 discusses them. Finally, Section 7 offers concluding remarks.
@@ 2 Related literature
Alternative indicators of long-term competitive balance have also been suggested. Groot, (2008, Chapter 1.4) proposes Kendall rank correlation to obtain a dynamic measure of competitive balance. Manasis et al., (2011) consider the normalised concentration ratio to evaluate the extent to which a league is dominated by a small subset of teams. Manasis et al., (2013) introduce an index to capture the degree of competition for winning the important prizes awarded in European football leagues such as the championship, the qualification for prestigious international club competitions, and avoiding relegation. Manasis and Ntzoufras, (2014) examine the applicability of between-season competitive balance indices in the context of European football, and define a set of specially designed measures. Hood and Jewell, (2022) illustrate the value of using betting data to simulate an ex-ante distribution of league-point outcomes in English football. Basini et al., (2023) develop a statistical network model to assess the balance between teams in a league. Csató and Petróczy, (2024) adopt bibliometric indices for quantifying competitive balance in the UEFA Champions League knockout stage. A recent empirical analysis of nontransitive patterns in professional football (van Ours,, 2024) can provide a novel approach to measuring competitive balance, too.
Traditional measures of performance have recently been challenged in multi-stage cycling races (Ausloos, 2024a, ; Ausloos, 2024b, ). Our paper has a similar aim in quantifying competitive balance in the group stage of (football) tournaments.
@@ 3 Potential limitations of previous studies
Triguero-Ruiz and Avila-Cano, (2023) start to quantify ex ante competitive balance by calculating the shares of the clubs from UEFA club coefficients. They are used to compute for each group, which gives the distance to competitive balance () index (Dubois,, 2022; Scelles et al.,, 2022; Triguero Ruiz and Avila-Cano,, 2019) as follows:
UEFA club coefficient is the official measure of team strength in UEFA club competitions, the number of points gained by a team in UEFA club competitions over the last five years (Dagaev and Rudyak,, 2019). There is a lower bound based on the five-season association coefficient, which might be relevant for emerging clubs without meaningful experience at the international level (Csató,, 2022). Primarily, the rankings derived from the UEFA club coefficients have determined seeding in the UEFA club competitions (Csató,, 2020).
@@
Ex post competitive balance is usually quantified by the Herfindahl–Hirschman index () based on the shares of each team from the total points awarded (Avila-Cano et al.,, 2023; Owen and Owen,, 2022; Owen et al.,, 2007). Both Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023) use this approach to compute ex post competitive balance in a group. The underlying assumption is that the teams have appropriate incentives to collect points even though their qualification depends on the group ranking rather than on the number of points. Therefore, the clubs do not necessarily have appropriate incentives in the last round(s) (Chater et al.,, 2021; Csató et al.,, 2024). Buraimo et al., (2022) find robust evidence that matches of no importance for seasonal outcomes—which is equivalent to group ranking in our setting—may not be important to viewers in the English Premier League.
@@
Based on the arguments above, two indices of ex ante and four indices of ex post competitive balance will be constructed in order to check the robustness of previous findings by analysing alternative but reasonable indicators of competitive balance. Naturally, our measures have their own weaknesses, and they should not necessarily be preferred to the indicators used in the previous literature.
@@ 4 Methodology and data
Sections 4.1 and 4.2 motivate and define the indices of ex ante and ex post competitive balance, respectively. Section 4.3 provides an example of their calculation, Section 4.4 gives information about the underlying data, while Section 4.5 presents how the dynamics of competitive balance is analysed. Last but not least, Section 4.6 summarises the main steps of our research.
@@ 4.1 Measures of ex ante competitive balance
Following Bosker and Gürtler, (2024) and Csató, (2024), we use Football Club Elo Ratings as a measure of team strength (Football Club Elo Ratings,, 2024). It gives the expected probability that team wins against team as:
@@
4.2 Measures of ex post competitive balance
We compute the Kendall rank correlation coefficient (Kendall,, 1938) between an ex ante ranking and the final group ranking. The use of Kendall rank correlation for competitive balance goes back at least to Groot, (2008), who applied it to evaluate dynamic competitive balance in football leagues. In a sense, ex post competitive balance in the UEFA Champions League groups is about mobility: an ex ante (before the season) ranking of the teams can be compared to the final (after the season) group ranking to determine the uncertainty of tournament outcome. Mobility is maximal if the final ranking is exactly the reverse of the ex ante ranking, corresponding to a perfect negative rank correlation. However, then one can perfectly predict the final ranking by reversing the initial order of the teams. Hence, the ideal situation is if the final ranking is random, that is, Kendall rank correlation equals zero (Groot,, 2008).
Denote the number of concordant and discordant pairs between the ex ante and ex post rankings by and , respectively, and the number of teams by . The Kendall rank correlation is
@@ 4.4 Data
We consider all seasons between 2003/04 and 2023/24 when the tournament format of the UEFA Champions League did not changed. The pot allocation and group rankings are readily available from several sources; we have used Wikipedia after cross-checking with the official UEFA site.
@@ 4.5 The evolution of competitive balance
The approach of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023) is followed directly. For all the six (two ex ante and four ex post) indicators of competitive balance, the linear trend is obtained by estimating the following regression model:
where is the th competitive balance measure and is the error term in season (). Every season is denoted by its first year when the group stage is played. Furthermore, is the intercept and is the coefficient of the season. Competitive balance declined/did not changed/improved if is significantly positive/statistically zero/significantly negative.
@@ 5 Results
Notes: The seasons are denoted by their first year when the group stage is played.
@@
Figure 1 shows the evolution and the linear trend of the six competitive balance indices. The means of ex post competitive balance measures are around 0.5, and even the smallest seasonal average is above 0.2. Therefore, both pot allocation and Elo ranking are better predictors of the final group ranking than a random permutation (Groot,, 2008) as expected. Group G in the 2019/20 UEFA Champions League is an interesting outlier, where the final ranking is the opposite of the ranking implied by the pot allocation.
@@
Figure 1 may suggest a small drop in competitive balance over the recent decades since the slope of the linear trend is always positive. However, none of them are significant according to Table 2. This is in stark contrast with Triguero-Ruiz and Avila-Cano, (2023), where the -values are below 0.02 for ex ante and approximately 0 for ex post competitive balance, respectively. Unsurprisingly, never exceeds 0.1, contrary to the results of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023).
@@ 6 Discussion
Regarding ex post competitive balance, we compare the final group ranking to a benchmark provided by either the pot allocation or the Elo ranking. Since the does not require such a benchmark, a worsening ex post competitive balance may be covered if our benchmark becomes a more accurate predictor over time. However, this is improbable. Even though the seeding system has been reformed as discussed in Section 3, it has rather increased uncertainty (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). Similarly, the differences in the Elo ratings have not increased (Figure 1), thus, the quality of Elo ranking seems to be unchanged. On the other hand, the clubs may focus in the group stage more strongly on their number of points than previously, possibly due to the increasing financial incentives provided by the UEFA. This leads to a higher and a lower probability of losing stakeless games. Then the earlier studies have only mixed a favourable trend (teams do not lose their stakeless matches) with declining ex post competitive balance. The hypothesis may be tested by an empirical investigation of stakeless games, which can be identified, for example, by the approach of Csató et al., (2024). Finally, it should be noted that the sample size is rather small. However, the novel format of the UEFA Champions League from the 2024/25 season does not allow for more observations.
@@ 7 Conclusions
UEFA has chosen a risky strategy by fundamentally reforming the well-established group stage of its major tournament, the UEFA Champions League, from the 2024/25 season. According to the existing literature based on straightforward -based measures, UEFA should have acted effectively because of the notable drop in both the ex ante and ex post competitive balance of the group stage over the last two decades. Our paper has proposed six alternative but reasonable indices to see the robustness of these results. Crucially, no evidence is found for a declining competitive balance between the 2003/04 and 2023/24 seasons. Therefore, we urge researchers and analysts to consider a broader set of competitive balance indicators in the future before proposing powerful policy interventions.
@@ Acknowledgements
This paper could not have been written without Réka Boros and Adrienn Czakó, who prepared the data and made the calculations.
@@
One anonymous reviewer provided useful comments on an earlier draft.
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838 and PD 146055, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.

### v6

# Competitive balance in the UEFA Champions League group stage: Novel measures show no evidence of decline

@@ Abstract
Competitive balance, which refers to the level of control teams have over a sports competition, is a crucial indicator for tournament organisers. According to previous studies, competitive balance has significantly declined in the UEFA Champions League group stage over the recent decades. Our paper introduces alternative indices to investigate this issue. Two ex ante measures are based on Elo ratings, and four dynamic concentration indicators compare the final group ranking to reasonable benchmarks. Using these indices, we find no evidence of any long-run trend in the competitive balance of the UEFA Champions League group stage between the 2003/04 and 2023/24 seasons.
@@ 1 Introduction
One of the most prestigious football tournaments, the UEFA Champions League, has been organised in the same format over 21 years, which has been fundamentally changed from the 2024/25 season. In particular, the traditional first stage, where four teams have played a home-away round-robin tournament in eight independent groups, has been replaced by a single incomplete round-robin league (Devriesere et al.,, 2025). Commenting on this decision, the president of the Union of European Football Associations (UEFA), Aleksander Čeferin, said (UEFA,, 2022): “We are convinced that the format chosen strikes the right balance and that it will improve the competitive balance and generate solid revenues that can be distributed to clubs, leagues and into grassroots football across our continent while increasing the appeal and popularity of our club competitions.” Indeed, the new competition design can be successful in improving competitive balance as shown by Gyimesi, (2024).
Consequently, this reform might have been inspired by problems perceived around the competitive balance of the UEFA Champions League. Recent results from academic literature support this assumption: a robust decline has been observed in both the ex ante (Triguero-Ruiz and Avila-Cano,, 2023) and ex post (Ramchandani et al.,, 2023; Triguero-Ruiz and Avila-Cano,, 2023) competitive balance of the UEFA Champions League group stage over the last decades.
Nonetheless, competitive balance is widely recognised to be a multi-dimensional concept (Sanderson,, 2002; Kringstad and Gerrard,, 2007; Gerrard and Kringstad,, 2022), and there exist several measures for it (Doria and Nalebuff,, 2021; Gerrard and Kringstad,, 2023; Manasis et al.,, 2022; Pawlowski and Nalbantis,, 2019). Our paper introduces and explains two alternative approaches to quantify competitive balance in the UEFA Champions League group stage.
The first is a standard ex ante competitive balance measure, which evaluates the strengths of the participating teams based on past performance. While UEFA has an official indicator of team strength called UEFA club coefficient, used for seeding in the UEFA Champions League (Csató,, 2020), there is a more accurate measure of team strength called Football Club Elo Ratings (Csató,, 2024). Using Elo ratings instead of UEFA club coefficients has the additional advantage that our ex ante competitive balance measure will have a probabilistic interpretation.
The second indicator, called dynamic concentration, is based on the idea of Groot, (2008) to measure the ranking turnover between two seasons of a league by Kendall’s coefficient. This method is adopted to compare an ex ante ranking to the ex post ranking in the groups within the same season. Hence, our measure essentially applies a standard interpretation of between-season competitive balance to one season. The motivation is that the rules of the UEFA Champions League group stage have not created powerful incentives to increase the number of points once the final rank of a team has become fixed (Csató et al.,, 2024; Guyon,, 2022).
@@
Based on these indices, we find no evidence of any decline in the competitive balance of the UEFA Champions League group stage between the 2003/04 and 2023/24 seasons. Therefore, if UEFA has chosen the barely used incomplete round-robin design of the UEFA Champions League with its inherent challenges and risks (Csató et al.,, 2025) because of the worsening trend in competitive balance, the decision-makers might have been misled.
Naturally, the proposed indicators may have their own disadvantages and shortcomings, and we do not claim that they should be preferred to existing indices. However, it would be unreasonable to dismiss them for the sake of the standard measures of competitive balance as “there are no easy formulas for competitive balance success” (Zimbalist,, 2002, p. 120).
@@
We think it is the responsibility of the academic community to consider as many measures as possible, which will be done in the following structure. Section 2 provides a concise literature review on measuring competitive balance. Section 3 summarises some possible limitations of previous studies. Our measures of competitive balance are introduced in Section 4 together with a description of the underlying data. Section 5 provides the results and Section 6 discusses them. Finally, Section 7 offers concluding remarks.
@@ 2 Related literature
Alternative indicators of long-term competitive balance have also been suggested. Groot, (2008, Chapter 1.4) proposes Kendall rank correlation to obtain a dynamic measure of competitive balance. Manasis et al., (2011) consider the normalised concentration ratio to evaluate the extent to which a league is dominated by a small subset of teams. Manasis et al., (2013) introduce an index to capture the degree of competition for winning the important prizes awarded in European football leagues such as the championship, the qualification for prestigious international club competitions, and avoiding relegation. Manasis and Ntzoufras, (2014) examine the applicability of between-season competitive balance indices in the context of European football, and define a set of specially designed measures. Hood and Jewell, (2022) illustrate the value of using betting data to simulate an ex-ante distribution of league-point outcomes in English football. Basini et al., (2023) develop a statistical network model to assess the balance between teams in a league. Gerrard and Kringstad, (2023) extend previous empirical studies on the four North American major sports leagues using six competitive balance measures. Similar to us, they conclude that the evaluation of competitive balance is both metric-dependent and time-dependent, which highlights the importance of considering a portfolio of indices rather than a single indicator. Csató and Petróczy, (2024) adopt bibliometric indices for quantifying competitive balance in the UEFA Champions League knockout stage. A recent empirical analysis of nontransitive patterns in professional football (van Ours,, 2024) can provide a novel approach to measuring competitive balance, too.
Traditional measures of performance have recently been challenged in multi-stage cycling races (Ausloos, 2024a, ; Ausloos, 2024b, ). Our paper has a similar aim by suggesting novel competitive balance measures in the group stage of (football) tournaments.
@@ 3 Potential limitations of previous studies
Compared to national leagues, measuring competitive balance in the UEFA Champions League has an additional challenge, that is, a team participating in a given year may not compete in the following year but replaced by a stronger team from the same domestic league (Koning,, 2009). Therefore, the strength of the teams should be quantified independently of their identity.
Triguero-Ruiz and Avila-Cano, (2023) start to measure ex ante competitive balance by calculating the shares of the clubs from UEFA club coefficients. They are used to compute for each group, which gives the distance to competitive balance () index (Dubois,, 2022; Scelles et al.,, 2022; Triguero Ruiz and Avila-Cano,, 2019) as follows:
UEFA club coefficient is the official measure of team strength in UEFA club competitions, the number of points gained by a team in UEFA club competitions over the last five years (Dagaev and Rudyak,, 2019). There is a lower bound based on the five-season association coefficient, which might be relevant for emerging clubs without meaningful experience at the international level (Csató,, 2022). Primarily, the rankings derived from the UEFA club coefficients have determined seeding in the UEFA club competitions (Csató,, 2020; Csató and Ilyin,, 2025).
@@
Ex post competitive balance is usually quantified by the Herfindahl–Hirschman index () based on the shares of each team from the total points awarded (Avila-Cano et al.,, 2023; Owen and Owen,, 2022; Owen et al.,, 2007). Both Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023) use this approach to compute ex post competitive balance in a group. The underlying assumption is that the teams have appropriate incentives to collect points even though their qualification depends on the group ranking rather than on the number of points. Thus, the clubs do not necessarily have appropriate incentives in the last round(s) (Chater et al.,, 2021; Csató et al.,, 2024). Buraimo et al., (2022) find robust evidence that matches of no importance for seasonal outcomes—which is equivalent to group ranking in our setting—may not be important to viewers in the English Premier League.
@@
Nevertheless, in contrast to the distribution of the results obtained, the final group ranking cannot be evaluated in itself without any benchmark. In other words, the group ranking could not serve as an ex post competitive balance measure because it should be compared to an ex ante ranking.
Based on the arguments above, two indices of ex ante competitive balance and four indices of dynamic concentration will be constructed in the following. Naturally, our measures have their own weaknesses, and they should not necessarily be preferred to the indicators used in the previous literature.
@@ 4 Methodology and data
Sections 4.1 and 4.2 motivate and define the indices of ex ante competitive balance and dynamic concentration, respectively. Section 4.3 provides an example of their calculation, Section 4.4 gives information about the underlying data, while Section 4.5 presents how the dynamics of competitive balance is analysed. Last but not least, Section 4.6 summarises the main steps of our research.
@@ 4.1 Measures of ex ante competitive balance
Following the recent literature (Bosker and Gürtler,, 2024; Csató,, 2024; Yildirim and Bilman, 2025a, ; Yildirim and Bilman, 2025b, ), we use Football Club Elo Ratings as a measure of team strength (Football Club Elo Ratings,, 2024). It gives the expected probability that team wins against team as:
@@
4.2 Measures of dynamic concentration
We compute the Kendall rank correlation coefficient (Kendall,, 1938) between an ex ante ranking and the final group ranking. The use of Kendall rank correlation for competitive balance goes back at least to Groot, (2008), who applied it to evaluate dynamic competitive balance in football leagues. In a sense, competitive balance in the UEFA Champions League groups is about mobility: an ex ante (before the season) ranking of the teams can be compared to the final (after the season) group ranking to determine the uncertainty of tournament outcome. Mobility is maximal and dynamic concentration is minimal if the final ranking is exactly the reverse of the ex ante ranking, corresponding to a perfect negative rank correlation. However, then one can perfectly predict the final ranking by reversing the initial order of the teams. Hence, the ideal situation is if the final ranking is random, that is, Kendall rank correlation equals zero (Groot,, 2008).
Denote the number of concordant and discordant pairs between the ex ante and ex post rankings by and , respectively, and the number of teams by . The Kendall rank correlation equals
@@ 4.4 Data
We consider all seasons between 2003/04 and 2023/24 when the tournament format of the UEFA Champions League did not change. The pot allocation and group rankings are readily available from several sources; we have used Wikipedia after cross-checking with the official UEFA site.
@@ 4.5 The evolution of competitive balance
The approach of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023) is followed directly. For all the six indicators of competitive balance, the linear trend is obtained by estimating a simple regression model:
where is the th competitive balance measure and is the error term in season (). Every season is denoted by its first year when the group stage is played. Furthermore, is the intercept and is the coefficient of the season. Competitive balance declined/did not change/improved if is significantly positive/statistically zero/significantly negative.
@@ 5 Results
Notes: The seasons are indicated by their first year when the group stage is played.
@@
Figure 1 shows the evolution and the linear trend of the six competitive balance indices. The means of dynamic concentration measures are around 0.5, and even the smallest seasonal average is above 0.2. Therefore, both pot allocation and Elo ranking are better predictors of the final group ranking than a random permutation (Groot,, 2008) as expected. Group G in the 2019/20 UEFA Champions League is an interesting outlier, where the final ranking was the opposite of the ranking implied by the pot allocation.
@@
Figure 1 may suggest a small drop in competitive balance over the recent decades since the slope of the linear trend is always positive. However, none of them are significant according to Table 2. This is in stark contrast with Triguero-Ruiz and Avila-Cano, (2023), where the -values are below 0.02 for ex ante competitive balance and approximately 0 for dynamic concentration, respectively. Unsurprisingly, never exceeds 0.1, contrary to the results of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023).
@@ 6 Discussion
Regarding dynamic concentration, we compare the final group ranking to a benchmark provided by either the pot allocation or the Elo ranking. Since does not require such a benchmark, a worsening competitive balance may be covered if our benchmark becomes a more accurate predictor over time. However, this is improbable. Even though the seeding system has been reformed as discussed in Section 3, it has rather increased uncertainty (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). Similarly, the differences in the Elo ratings have not increased (Figure 1), thus, the quality of Elo ranking seems to be unchanged. On the other hand, the clubs may focus in the group stage more strongly on their number of points than previously, possibly due to the increasing financial incentives provided by the UEFA. This could lead to a higher and a lower probability of losing stakeless games. Then the earlier studies have mixed a favourable trend (teams do not lose their stakeless matches) with declining ex post competitive balance. The hypothesis may be tested by an empirical investigation of stakeless games, which can be identified, for example, by the approach of Csató et al., (2024). Finally, it should be noted that the sample size is rather small. However, the novel format of the UEFA Champions League from the 2024/25 season does not allow for more observations.
@@ 7 Conclusions
UEFA has chosen a risky strategy by fundamentally reforming the well-established group stage of its major tournament, the UEFA Champions League, from the 2024/25 season. According to the existing literature based on straightforward -based measures, UEFA should have acted effectively because of the notable drop in both the ex ante and ex post competitive balance of the group stage over the last two decades.
Our paper has proposed six alternative but reasonable indices to investigate the evolution of competitive balance in the UEFA Champions League group stage. Crucially, no evidence is found for any decline between the 2003/04 and 2023/24 seasons. Therefore, we urge researchers and analysts to consider a broader set of competitive balance indicators in the future before proposing powerful policy interventions.
@@ Acknowledgements
This paper could not have been written without Réka Boros and Adrienn Czakó, who prepared the data.
@@
One anonymous reviewer provided valuable remarks on earlier drafts.
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838 and PD 153835, and by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.